Differentiate the following functions:



1.y=1/2(3x^2+1)^2



2.y=3+4x-x^2



Find the derivative of the following



1.y=√x^2-2/2x^3+1



2.y=x^2√6x^2/3x



3.y=2x^2√x

1
Expert's answer
2022-03-22T18:35:24-0400

"1) y=\frac{1}{2(3x^2+1)^2}\\
y'=\frac{1}{2}(-2)\frac{1}{(3x^2+1)^3}\cdot(6x)=\\
=\frac{-6x}{(3x^2+1)^2}\\
2) y=3+4x-x^2\\
y'=0+4-2x=4-2x\\
3) y=\frac{\sqrt{x^2-2}}{2x^3+1}\\
y'=\frac{\frac{1}{2\sqrt{x^2-2}}\cdot 2x\cdot (2x^3+1)- \sqrt{x^2-2}\cdot 6x^2}{(2x^3+1)^2}=\\
=\frac{x\cdot (2x^3+1)- (x^2-2)\cdot 6x^2}{\sqrt{x^2-2}(2x^3+1)^2}=\\
=\frac{-4x^4+12x^2+x}{\sqrt{x^2-2}(2x^3+1)^2}\\
4) y=\frac{x^2\sqrt{6x^2}}{3x}\\
y=\frac{x\sqrt{6x^2}}{3}=\frac{\sqrt{6}}{3}x^2\\
y'=\frac{\sqrt{6}}{3}2x=\frac{2\sqrt{6}x}{3}\\
5) y=2x^2\sqrt{x}=2x^{\frac{5}{2}}\\
y'=2\cdot\frac{5}{2}\cdot x^{\frac{3}{2}}=5x^{\frac{3}{2}}"